### v1

@@ Abstract
Recent advances in multimodal models have demonstrated impressive capabilities in object recognition and scene understanding. However, these models often struggle with precise spatial localization - a critical capability for real-world applications. Inspired by how humans use grid-based references like chess boards and maps, we propose introducing explicit visual position encoding through a simple grid overlay approach. By adding a 9×9 black grid pattern onto input images, our method provides visual spatial guidance analogous to how positional encoding works in transformers, but in an explicit, visual form.
Experiments on the COCO 2017 dataset demonstrate that our grid-based approach achieves significant improvements in localization accuracy, with a 107.4% increase in IoU (from 0.27 to 0.56) and a 194.4% improvement in GIoU (from 0.18 to 0.53) compared to baseline performance. Through attention visualization analysis, we show how this visual position encoding helps models better ground spatial relationships. Our method’s simplicity and effectiveness make it particularly valuable for applications requiring accurate spatial reasoning, such as robotic manipulation, medical imaging, and autonomous navigation.
@@ 1 Introduction
Our work is motivated by a fundamental question: Can we make positional encoding more explicit and visual? Drawing inspiration from how humans use grid-based references (like chess boards or map coordinates) for precise spatial communication, we propose a novel approach to spatial understanding in multimodal models. Instead of relying on implicit numerical positional encodings, we introduce a visual grid system that serves as an explicit, visual form of positional encoding. By overlaying a grid pattern onto input images, we provide direct spatial reference points that the model can use for precise localization.
@@ 2.1 Position Information in Vision-Language Models
Modern vision-language models have revolutionized our ability to understand and reason about visual content.[12] At their core, these models must address a fundamental challenge: how to effectively encode and utilize spatial information. We begin by examining how different approaches handle this crucial aspect.
@@ 2.2.1 Traditional Vision Backbone Approaches
$f_{spatial}(x)=CNN(x)\rightarrow abstract\>features$ (3)
@@ 2.2.3 Cross-Modal Spatial Understanding
$p(location|text,image)=softmax(f_{cross-attention}(text,image\>features))$ (5)
@@ 2.2.4 Direct Coordinate Prediction Approaches
$(x,y)=f_{regression}(image\>features)$ (6)
However, these approaches often suffer from training instability and scale sensitivity. More sophisticated approaches like DETR [14] use learned object queries and direct set prediction of bounding boxes, enabling end-to-end training without traditional positional encoding. While these methods show promise, they still face significant challenges in convergence speed, small object detection, and achieving the level of precision required for exact coordinate prediction in practical applications.
@@ 2.3.2 Grid-Based References in Computer Vision
$location=f_{grid}(row,column)\rightarrow explicit\>spatial\>reference$ (8)
@@ 2.4 Prompt Engineering and Visual Guidance
The emergence of prompt engineering in vision-language models offers interesting parallels to our work. Similar to how textual prompts guide language models, visual prompts can provide explicit guidance for spatial understanding. SAM [16] demonstrated the power of visual prompts for segmentation tasks, though it focuses on region selection rather than coordinate prediction.
@@ 4.4 Visualization Results
When tested on the original image without grid overlay, the model’s predictions exhibited substantial inaccuracies, with predicted coordinates deviating significantly from the ground truth location. However, after applying our $9\times 9$ grid reference system, we observed a marked improvement in localization accuracy. The grid’s presence provided explicit spatial anchors that enabled the model to generate more precise coordinate predictions, effectively reducing the localization error. This improvement demonstrates how our simple grid-based approach can enhance the model’s ability to perform accurate spatial localization.
@@ 5 Discussion
The effectiveness of our grid-based approach provides several key insights into spatial reference mechanisms in neural networks. While traditional backbone architectures excel at feature extraction, our results suggest that explicit spatial reference information can serve as a vital complement for precise localization tasks. The grid overlay effectively acts as spatial guidances, providing the model with consistent reference points that enhance its ability to make accurate coordinate predictions. Notably, this improvement does not rely on complex architectural modifications but is achieved through the provision of clear spatial guidance. This finding challenges the conventional wisdom that increasing model complexity is a necessary path to improving spatial understanding.

### v2

@@ Abstract
Recent advances in multimodal models have demonstrated impressive capabilities in object recognition and scene understanding. However, these models often struggle with precise spatial localization - a critical capability for real-world applications. We propose a simple yet effective backbone-free approach that introduces an explicit visual reference system for enhanced spatial understanding. Byoverlaying a 9×9 black grid pattern onto input images, our method provides consistent spatial anchors without requiring architectural modifications or additional computational overhead. Experiments on the COCO 2017 dataset demonstrate that our grid-based approach achieves significant improvements in localization accuracy, with a 107.4% increase in IoU (from 0.27 to 0.56) and a 194.4% improvement in GIoU (from 0.18 to 0.53) compared to baseline performance. These results highlight how explicit spatial references can effectively bridge the gap between conceptual understanding and precise localization. Our method’s simplicity and effectiveness make it particularly valuable for applications requiring accurate spatial reasoning, such as robotic manipulation, medical imaging, and autonomous navigation. The project code is available at https://github.com/triumph123aaa/GRID-AUGMENTED-VISION .
@@ 1 Introduction
Our work is motivated by a fundamental question: Can we make positional encoding more explicit and visual? Drawing inspiration from how humans use grid-based references (like chess boards or map coordinates) for precise spatial communication, we propose a novel approach to spatial understanding in multimodal models. Instead of relying on implicit numerical positional encodings, we introduce a visual grid system that serves as an explicit, visual form of positional encoding. By overlaying a 9 × 9 black grid pattern onto input images, we provide direct spatial reference points that the model can use for precise localization.
@@ 2.1 Position Information in Vision-Language Models
Modern vision-language models have revolutionized our ability to understand and reason about visual content. At their core, these models must address a fundamental challenge: how to effectively encode and utilize spatial information. We begin by examining how different approaches handle this crucial aspect.
@@ 2.2.1 Traditional Vision Backbone Approaches
$f_{spatial}(x)=CNN(x)\rightarrow abstract\ features$ (3)
@@ 2.2.3 Cross-Modal Spatial Understanding
$p(location|text,image)=softmax(f_{cross-attention}(text,image\ features))$ (5)
@@ 2.2.4 Direct Coordinate Prediction Approaches
$(x,y)=f_{regression}(image\ features)$ (6)
However, these approaches often suffer from training instability and scale sensitivity. More sophisticated approaches like DETR [13] use learned object queries and direct set prediction of bounding boxes, enabling end-to-end training without traditional anchor boxes. While these methods show promise, they still face significant challenges in convergence speed, small object detection, and achieving the level of precision required for exact coordinate prediction in practical applications.
@@ 2.3.2 Grid-Based References in Computer Vision
$location=f_{grid}(row,column)\rightarrow explicit\ spatial\ reference$ (8)
@@ 2.4 Prompt Engineering and Visual Guidance
The emergence of prompt engineering in vision-language models offers interesting parallels to our work. Prompts, functioning as instructional directives or conditional parameters supplied to artificial intelligence models, can manifest in either textual or visual modalities. Similar to how textual prompts guide language models, visual prompts can provide explicit guidance for spatial understanding. SAM [15] demonstrated the power of visual prompts for segmentation tasks, though it focuses on region selection rather than coordinate prediction.
@@ 4.4 Visualization Results
When tested on the original image without grid overlay, the model’s predictions exhibited substantial inaccuracies, with predicted coordinates deviating significantly from the ground truth location. However, after applying our $9\times 9$ grid reference system, we observed a marked improvement in localization accuracy.We conducted multiple comparative experiments between original images and grid-overlaid images, with detailed results presented in Appendix  A. The grid’s presence provided explicit spatial anchors that enabled the model to generate more precise coordinate predictions, effectively reducing the localization error. This improvement demonstrates how our simple grid-based approach can enhance the model’s ability to perform accurate spatial localization.
@@ 5 Discussion
The effectiveness of our grid-based approach provides several key insights into spatial reference mechanisms in neural networks. While traditional backbone architectures excel at feature extraction, our results suggest that explicit spatial reference information can serve as a vital complement for precise localization tasks. The grid overlay effectively acts as spatial anchors, providing the model with consistent reference points that enhance its ability to make accurate coordinate predictions. Notably, this improvement does not rely on complex architectural modifications but is achieved through the provision of clear spatial guidance. This finding challenges the conventional wisdom that increasing model complexity is a necessary path to improving spatial understanding.
@@ 6 Conclusion
Appendix A More Results
In our experiments, we employed the following prompt to guide the model in person detection and bounding box localization tasks: Please find the bounding box coordinates of the person in this image. When scanning the image to locate the person, look across the entire image and consider all visible human figures. Once you have found the person, determine their exact boundaries by identifying the left, right, top, and bottom boundaries. Please provide your reasoning process and give the final coordinates as [x1, y1, x2, y2] for the bounding box. To comprehensively validate the effectiveness of our methodology, we conducted comparative experiments across 9 representative scenarios. Each comparison pair comprises predictions from the original image (left) and its corresponding version with a 9×9 grid overlay (right), where we visualized the bounding box coordinates predicted by the model. As illustrated in Figure 5, the comparative results clearly demonstrate the enhanced person localization accuracy after introducing the grid structure, with the left image showing bounding box predictions on the original image and the right image displaying predictions with the 9×9 reference grid overlay. The experimental results indicate that the incorporation of the grid structure significantly improved the model’s precision in person bounding box localization.
Figure 5: Visualization of person localization results across nine different scenarios. For each pair, the left image shows predictions on the original image, while the right image displays predictions with a 9×9 grid overlay, demonstrating improved localization accuracy with the grid-based approach.